Solve the following problems on Price elasticity of demand and classify the nature of elasticity and good. 1) The demand for a CDs is 1000 units and the price is Rs. 20/- per CD. Later on the price rises to Rs.22/- per CD as result the demand falls to 870 units
1
Expert's answer
2021-10-28T16:23:50-0400

1.). Price elasticity of demand (PED) "=\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; price}"


PED = "=\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 }"


Where: Q1 = 1000 P1 = 20

Q2 = 870 P2 = 22


PED = "\frac{870 -1000}{(870+1000)/2 } \div\frac{22 -20}{(22+20)/2 }"


"\frac{-130}{935} \div\frac{2}{21} = \frac{-0.139}{0.095} = -1.46"


PED = 1.46

The good is price elastic since PED is greater than 1, indicating that the quantity demanded is highly responsive to price variations. The good is ordinary good.